How do I convert the decimal notation 0.23 to a percent notation?
Found 3 solutions by jim_thompson5910, richwmiller, Maths68:
Answer by jim_thompson5910(35256)

Multiply by 100 to get 0.23*100 = 23%

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------

Answer by richwmiller(17219)

.23=23/100=23 per hundred =23 per cent

Answer by Maths68(1474)

Remember
"per" means "out of"
"cent" means " hundred"
so
"percent" means "out of hundred"
Now come to your question.

=0.23
=23/100 (23 out of hundred)
=23%